OpenStudy (mathlegend):

Factor Quadratic Functions: Question.... Is there a way to factor quadratics in a fast way? How would you factor: 2t^2-20t+50=0

OpenStudy (anonymous):

Problem: Factor Quadratic Functions: Question.... Is there a way to factor quadratics in a fast way? How would you factor: 2t^2-20t+50=0 You're off to a good start. You made sure that the 0 is isolated. On the left side, now take out the common factor of the three terms. Know what that is? After doing that, factor what's left. Let me know if you're able to do that. Good luck!

OpenStudy (mathlegend):

common factor is 2

OpenStudy (mathlegend):

2(t^2-10t+25)=0 2(t-5)(t-5)=0

OpenStudy (mathlegend):

t= 5 t = 2 ??

OpenStudy (mathlegend):

@abcabc but that was too easy... I mean how could I solve something like 12v^2-25v-7

OpenStudy (anonymous):

do you know the box method... you take the factors of the first term, factors of the second, cross multiply to get the middle?? do you know that way

OpenStudy (anonymous):